OpenStudy (anonymous):

Topic: \(\textbf{Calculus 2}\) Q: Radius of convergence and interval of convergence for: \(\large \sum_{n=1}^{\infty} n!(2x-1)^n\) Ratio Test for Divergence \(\lim_{n \rightarrow \infty} \left| \frac{a_{n+1}}{a_n} \right|\) = L if ( L < 1 ) then {convergent} if( L > 1 ) then {"divergent"} else {"inconclusive"}

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \left| \frac{a_{n+1}}{a_n} \right|\] \[\lim_{n \rightarrow \infty} \left| \frac{(n+1)!(2x-1)^{n+1}}{n!(2x+1)^n} \right|\] \[\lim_{n \rightarrow \infty} \left| \frac{\cancel{n!}(n+1)(2x-1)^{\cancel{n}+1}}{\cancel{n!}\cancel{(2x+1)^n}} \right| = \infty\]

OpenStudy (anonymous):

(trying to copy over from the related discussion about the trick shown in the last step to get rid of the factorial)

OpenStudy (anonymous):

@agentx5, it looks right to me, if you are only asking for a double check.

OpenStudy (anonymous):

Interval of convergence I have correct, -0.5 < x < 0.5 by solving the conjunction" |2x-1|<0

OpenStudy (anonymous):

But the radius? Why is it zero? (correct answer) I would have said the radius is either the same as the interval or all real numbers...

OpenStudy (anonymous):

See what I mean @Spacelimbus ? This is the second time I've seen this type of problem, so there's some concept here about radius of convergence I'm not understanding I guess...

OpenStudy (anonymous):

I made an educated guess on zero after 1/9 and \(\infty\) failed to be correct

OpenStudy (anonymous):

Graphing this on a TI-83 shows it variates, it's not constant at zero.

OpenStudy (anonymous):

well what makes sense to me in the example you just posted @agentx5 is that the term (9x-1)^n approaches it's limit faster then n! , but then I would have said that this is only the case if it it smaller than 1.

OpenStudy (experimentx):

man this is bugging me!!

OpenStudy (experimentx):

if x=1/n, the the sum converges!! but i haven't seen radius of interval in terms of n

OpenStudy (anonymous):

yes @experimentX, usually they neatly cancel out.

OpenStudy (experimentx):

yeah they do ... still, for any numerical value of x, the series clearly diverges no matther how small ... but for variable value 1/n ... this series converges. I haven't seen radius of convergence expressed interms of n either ... and can't show from from ratio test either

OpenStudy (experimentx):

1/9

OpenStudy (experimentx):

that's the only value for which the series converges and it converges to 0 .. there's no interval ... just a point

OpenStudy (anonymous):

I believe what @experimentX said would work for both answers @agentx5

OpenStudy (anonymous):

Oh bigger picture question to this then: What IS the radius of convergence, really?

OpenStudy (anonymous):

In your previous example. x=1/2 is the only point where the series converges. And that is one individual point, so no radius.

OpenStudy (anonymous):

Like in a practical or visual sense so I can understand it, what is it? Interval is the range of inputs in which it would converge

OpenStudy (experimentx):

it's a point

OpenStudy (anonymous):

A radius of convergence would mean, any value which is inside that radius (think of a circle) would make the sum converge, but in the examples you have posted, there is only one point that makes that happen, a point has no radius.

OpenStudy (anonymous):